# Compression of Higher Order Ambisonics with Multichannel RVQGAN

### v1

@@ Abstract
A multichannel extension to the RVQGAN neural coding method is proposed, and realized for data-driven compression of third-order Ambisonics audio. The input- and output layers of the generator and discriminator models are modified to accept multiple (16) channels without increasing the model bitrate. We also propose a loss function for accounting for spatial perception in immersive reproduction, and transfer learning from single-channel models. Listening test results with 7.1.4 immersive playback show that the proposed extension is suitable for coding scene-based, 16-channel Ambisonics content with good quality at 16 kbit/s.
@@ 3 Methods
The basic idea of the present multichannel architecture extension can be summarized as simply increasing the channel count of the first and last layer or RVQGAN convolutional layers to match the number of audio channels. In the case of 3rd-order Ambisonics material, this channel count is 16. With such a change, it is possible to keep the dimensionality, and therefore the compression efficiency of the model bottleneck. Besides straightforward architectural changes, this implies non-trivial changes to the typical loss functions (see next Section).
@@
We propose applying transfer learning to the multichannel model from previously trained mono model weights. In input and output layers, where the number of channels is increased, we replicate the original convolutional weights of the original channel to the multiple channels of the new model identically. Thus, all input and output channels of the multichannel model are processed with identical convolutional kernels at the start of the training, which then start to evolve as needed.
In practice, we construct the new model and copy the appropriate parameters from the state dictionary of the published Descript monaural model [3]. In addition to helping with RVQGAN training time consumption, transfer learning also helps if the fine-tuning stage is performed on less powerful hardware and smaller batch size. Fig. 1 illustrates the benefit of the proposed transfer learning in both faster conversion, and final error with reasonable amount of steps.
@@
Figure 1: Example comparison between proposed transfer learning utilizing pre-trained mono model weights, and vanilla random initialization. Plot shows the development of the reconstruction multi-scale mel validation loss [3].
@@
The model implementation was built on the Descript codebase [3] with the aforementioned multichannel extensions. The only other change was to omit the output Tanh-nonlinearity of the decoder. The training procedure remains largely the same as in the original work, while we utilize the transfer learning scheme proposed in the previous Section . We use weighting of 1.0 for the covariance loss. The other loss weightings were kept as is (15.0 for the multi-scale mel loss, 2.0 for the feature matching loss, 1.0 for the adversarial loss and 1.0, 0.25 for the codebook and commitment losses, respectively). To illustrate a simple fine tuning procedure results, we trained for a modest number of steps (400000) and batch size (24) of 5-second samples.
In order to evaluate and compare the proposed Ambisonics neural codec, a MUSHRA [35] listening test was carried out. For quality reference, we process the HOA reference signals with Opus channel mapping family 3 coding intended for HOA content [36], at 160 kbit/s rate, which is on the lower quality region of the Opus HOA, to demonstrate how conventional coding methods with 10x bitrate compare. In total, four listening conditions were compared:
@@
  • Opus’ Channel Family Mapping 3 (OPUS) compression at 160 kbit/s
  • Proposed neural ambisonics codec (PROP) at 16 kbit/s
@@ 4 Results and Discussion
As can be seen from Fig. 2, the proposed neural codec operating on 16 kbit/s is able to achieve ”good” quality on the MUSHRA scale, and outperform a traditional method with 10x less compression rate. While the traditional channel family mapping Opus method starts to break down at 160 kbit/s, and its quality would improve at higher rates, we opted to illustrate the quality potential of the neural methods rather than find the precise rate where the two methods would match.
@@
The audio material used here is from a truncated distribution of ambient scenes, and also limited in amount in order to illustrate a typical use case, as discussed in the Ambisonics-section earlier. However, the use of transfer learning from a generic model arguably retains some level of baseline generality. Also, the listening test conditions did enable critical evaluation of the sound quality similarly as in typical immersive audio testing. We also did not include in the tests any mixed presentations with additional audio elements [11], so that the HOA coding quality would be more forgiving.
Despite the promising results, the main caveats of the present method, and most neural codecs remain 1) the lack of scaling to different bitrates without retraining or model modification, and 2) the model computational complexity. For these, future work will hopefully provide solutions.

### v2

@@ Abstract
A multichannel extension to the RVQGAN neural coding method is proposed, and realized for data-driven compression of third-order Ambisonics audio. The input- and output layers of the generator and discriminator models are modified to accept multiple (16) channels without increasing the model bitrate. We also propose a loss function for accounting for spatial perception in immersive reproduction, and transfer learning from single-channel models. Listening test results with 7.1.4 immersive playback show that the proposed extension is suitable for coding scene-based, 16-channel Ambisonics content with good quality at 16 kbps.
@@ 3 Methods
The basic idea of the present multichannel architecture extension can be summarized as simply increasing the channel count of the first and last convolutional layer of RVQGAN to match the number of audio channels. In the case of 3rd-order Ambisonics material, this channel count is 16. With such a change, it is possible to keep the dimensionality, and therefore the compression efficiency of the model bottleneck. Besides straightforward architectural changes, this implies non-trivial changes to the typical loss functions (see next section).
@@
We propose applying transfer learning to the multichannel model from previously trained single-channel model weights. In input and output layers, where the number of channels is increased, we replicate the original convolutional weights of the original channel to the multiple channels of the new model identically. Thus, all input and output channels of the multichannel model are processed with identical convolutional kernels at the start of the training, which then evolve as needed.
In practice, we construct the new model and copy the appropriate parameters from the state dictionary of the published 16 kbps version of Descript monaural model [3]. In addition to helping with RVQGAN training time consumption, transfer learning also helps if the fine-tuning stage is performed on less powerful hardware and smaller batch size. Fig. 1 illustrates the benefit of the proposed transfer learning in both faster conversion, and final error with reasonable amount of steps.
@@
Figure 1: Example comparison between proposed transfer learning utilizing pre-trained single-channel model weights, and vanilla random initialization. Plot shows the development of the reconstruction multi-scale mel validation loss [3].
@@
The model implementation was built on the Descript codebase [3] with the aforementioned multichannel extensions to the architecture and loss calculation. The only other changes were to omit the output Tanh-nonlinearity, and trivial modifications to make the dependencies work with 16-channel audio. The training procedure remains largely the same as in the original work, while we utilize the transfer learning scheme proposed in the previous section. We use weighting of 1.0 for the covariance loss. The other loss weightings were kept as is (15.0 for the multi-scale mel loss, 2.0 for the feature matching loss, 1.0 for the adversarial loss and 1.0, 0.25 for the codebook and commitment losses, respectively). To illustrate a simple fine tuning procedure results, we trained for a modest number of steps (400000) and batch size (24) of 5-second samples.
In order to evaluate and compare the proposed Ambisonics neural codec, a MUSHRA [35] listening test was carried out. For quality reference, we utilize Opus channel mapping family 3 coding intended for HOA content [36], at 160 kbps rate, which is on the lower quality region of the Opus HOA, to demonstrate how conventional coding methods with 10x bitrate compare. In total, four listening conditions were compared:
@@
  • Opus’ Channel Family Mapping 3 (OPUS) compression at 160 kbps
  • Proposed neural ambisonics codec (PROP) at 16 kbps
@@ 4 Results and Discussion
As can be seen from Fig. 2, the proposed neural codec operating on 16 kbps is able to achieve ”good” quality on the MUSHRA scale, and outperform a traditional method with 10x less compression rate. While the traditional channel family mapping Opus method starts to break down at 160 kbps, and its quality would improve at higher rates, we opted to illustrate the quality potential of the neural methods rather than find the precise rate where the two methods would match.
@@
The audio material used here is from a truncated distribution of ambient scenes, and also limited in amount in order to illustrate a typical use case, as discussed in the Ambisonics-section earlier. However, the use of transfer learning from a generic model arguably retains some level of baseline generality. Also, the listening test conditions did enable critical evaluation of the sound quality similarly as in typical immersive audio testing. Although the present material was ambient background in nature, we did not test mixed presentations with additional foreground audio elements [11], which would have made the HOA coding quality less critical.
Although not explicitly tested here, the present modeling principle is applicable to multichannel audio formats other than 3rd-order Ambisonics (e.g. 5.1, 7.1, 7.1.4 etc.), as long as there is sufficient training data. Despite the promising results, the main caveats of the present method, and most neural codecs remain 1) the lack of scaling to different bitrates without retraining or model modification, and 2) the model computational complexity. For these, future work will hopefully provide solutions.